OpenStudy (anonymous):

csc^4 x-cot^4 x= (1+cos^2 x)/1-cos^2 x

OpenStudy (anonymous):

I need to verify this identity

OpenStudy (solomonzelman):

\[\csc^4 x-\cot^4 x= \frac{(1+\cos^2 x)}{1-\cos^2 x}\] @Jeffrospin, this?

OpenStudy (anonymous):

Yes! Just without the parentheses (my typo).

OpenStudy (solomonzelman):